Solve the initial value problem y"+3y'=0; y(0)=3, y'(0)=6
1
Expert's answer
2019-09-12T09:00:27-0400

Solve the initial value problem y"+3y'=0; y(0)=3, y'(0)=6.

The characteristic equation is

"k^2+3k=0 \\
k(k+3)=0 \implies k_1=0, k_2=-3."

We have two real distinct roots, then the general solution is

"y(t)=C_1e^{0t}+C_2e^{-3t}=C_1+C_2e^{-3t}."

Note that "y'(t)=-3C_2e^{-3t}."

Applying the initial conditions gives the following system of equations for the coefficients "C_1, C_2" :

"3=y(0)=C_1+C_2, \quad (1)\\
6=y'(0)=-3C_2 \implies C_2=-2."

From (1): "C_1=3-C_2=3-(-2)=5".

We get the solution "y(t)=5-2e^{-3t}."